plot the speed time graph from the table given below and calculate the distance covered by the particle during the time interval t=0 and t=0.4 sec. also find acceleration from the graph. choose appropriate scale
1
Expert's answer
2020-05-14T09:03:37-0400

Consider the following table:



The velocity vs. time graph will be like this:


The distance is the area below the incline:


"d=\frac{1}{2}vt=\frac{1}{2}\cdot1.6\cdot0.4=0.32\text{ m}."

The acceleration is the tangent of the angle between the time axis and the incline:


"a=\frac{v}{t}=\frac{1.6}{0.4}=4\text{ m/s}^2."